OpenStudy (anonymous):

Solve each System: -2x+4y+4z=24 2z=8 4x+y=11

OpenStudy (anonymous):

-2x+4y+4z=24 Move all terms not containing y to the right-hand side of the equation. 4y=2x-4z+24 Divide each term in the equation by 4. (4y)/(4)=(2x)/(4)-(4z)/(4)+(24)/(4) Simplify the left-hand side of the equation by canceling the common factors. y=(2x)/(4)-(4z)/(4)+(24)/(4) Simplify the right-hand side of the equation by simplifying each term. y=(x-2z+12)/(2)

OpenStudy (anonymous):

2z=8 Divide each term in the equation by 2. (2z)/(2)=(8)/(2) Simplify the left-hand side of the equation by canceling the common factors. z=(8)/(2) Simplify the right-hand side of the equation by simplifying each term. z=4

OpenStudy (anonymous):

4x+y=11 Since 4x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4x from both sides. y=-4x+11

OpenStudy (anonymous):

4x+y=11 Since y does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting y from both sides. 4x=-y+11 Divide each term in the equation by 4. (4x)/(4)=-(y)/(4)+(11)/(4) Simplify the left-hand side of the equation by canceling the common factors. x=-(y)/(4)+(11)/(4) Combine the numerators of all expressions that have common denominators. x=(-y+11)/(4)